The volume of a right circular cone obtained from a sector of a circle in which the radius is 26 cm and and the central angle is 138.5


1
Expert's answer
2021-06-24T09:24:42-0400

Length of the arc is

"l=\frac{\theta}{360 \degree }\times 2\pi r\\
\theta=138.5\degree\\
r=26cm\\
l=\frac{138.5\degree}{360 \degree }\times 2\times \frac{22}{7}\times 26cm\\
l=62.9cm\\"

Length of arc is the same as the circumference of base




"l=2\pi r_b\\
r_b=\frac{l}{2\pi}\\
r_b=\frac{62.9}{2\pi}\\
r_b=10cm\\"

Slant height, cone height and base radius form a right-angle triangle

"l_s=\sqrt{r_b^2+h^2}\\
l_s=r=26cm\\
h=\sqrt{26^2-10^2}\\
h=24cm\\"

Volume of cone is

"V=\frac{1}{3}\pi r_b^2h\\
V=\frac{1}{3}\times\frac{22}{7}\times{10^2}\times{24}\\
V=2514cm^3"